OpenStudy (anonymous):

find the value of k such that the equation has one real root x^2+12+k=0

OpenStudy (dumbcow):

to have 1 root it needs to in the form (x-r)^2 = 0 (x-r)^2 = x^2 -2xr +r^2 -2r = 12 --> r = -6 r^2 = k --> k = 36

OpenStudy (dumbcow):

you could also think of completing the square: take half of 12 and square it

OpenStudy (anonymous):

I did not think of doing that thanks